OpenStudy (anonymous):

A local basketball team is selling chocolate bars for $1 each. The team paid $.57 per bar plus $12.40 shipping. How many chocolate bars must be sold to make a profit of at least $300?

OpenStudy (anonymous):

the profit for sell one bar is 1-0.57=0.43,and constant cost is 12.40 let they sell x bar \[0.43x-12.40\ge 300\] \[x \ge 726.5\] so \[x \ge727\]

OpenStudy (anonymous):

thanks :)